OpenStudy (anonymous):

determine whether the sequence is increasing,decreasing or bounded, ln(n+2)/n+2

OpenStudy (anonymous):

For answering the question we consider the function : \[f(x)=\frac{\ln(x+2)}{x+2}\quad x\geq0.\] We derive it : \[f'(x)=\frac{\frac1{x+2}\times(x+2)-\ln(x+2)}{(x+2)^2}=\frac{1-\ln(x+2)}{(x+2)^2} \] And we have : \[1-\ln(x+2)=0\iff \ln(x+2)=1\iff x+2=e\iff x=e-2\simeq=0.7\] So we can sat that : \[\forall x\geq 1,\quad f'(x)<0\] We have : \[f(1)=\frac{\ln 3}3\\ \lim_{x\to+\infty}f(x)=0 \] So : \[1) ~~f~\text{ is decreasing over }[1,+\infty(\\ 2)~~\forall x\geq1,\quad 0\leq f(x)\leq \frac{\ln3}3. \] Now, our progression is defined by : \[u_n=\frac{\ln(n+2)}{n+2}=f(n)\] So : \[1) \left(u_n\right)_{n\in{\mathbb N}}\text{ is decreasing }\\ 2) \left(u_n\right)_{n\in{\mathbb N}}\text{ is bounded } \]

OpenStudy (anonymous):

Thanks a million.let me add for u,would u help me pls?

OpenStudy (anonymous):

using definition show that lim n/2n-3 = 1/2 as as n approches to infinity